OpenStudy (anonymous):

Let g(s)=(5s−4)8. Then g′(s) is________ and g′′(s) is _______ I dont know why I cant get the right answer for this question....

OpenStudy (anonymous):

g prime of (s) is 5(8) which = 40 and you cant take the derivative of 40

OpenStudy (anonymous):

well, i believe the derivative of a constant is 0. so g''(5s-4)8=0

OpenStudy (anonymous):

the first derivative is indeed 40. i hope this helps. :)

OpenStudy (anonymous):

oh wait I made a typo! Sorry guys! that is supposed to be to the power of 8 not x 8

OpenStudy (anonymous):

(5s-4)^8

OpenStudy (anonymous):

so if y=x^8, what is y'=?

OpenStudy (anonymous):

you have to use the chain rule for the problem

OpenStudy (anonymous):

i made a mistake in mine

OpenStudy (anonymous):

forgot your exponents

OpenStudy (anonymous):

ooooo and that's because the brackets counts as one part and the power as another, is that correct? Thats why the chain rule applies?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is why i was gonna start you with x^8

OpenStudy (anonymous):